4 (-4) / (-8) 6 = -3 + 5(3)
On the textbook it says that it is false

Answer by josgarithmetic(39617)

Left member connects its numbers only through multiplications and divisions.

4 (-4) / (-8) 6 = -3 + 5(3)
-16/(-8)*6=-3+15
2*6=12
12=12
The equation is TRUE.

Work strictly from left to right on expressions built only by multiplications and divisions.